The prelim2e package^{∗}

Martin Schröder

Friedrich-Humbert-Straße 124 D-28759 Bremen MS@Dream.HB.North.DE (INTERNET)

1996/01/01

Abstract

This package allows the marking of (preliminary) versions of a document on the output.

Contents

1 Introduction 1

2 Usage 1

3 Options 2 4 Required packages 2

5 Putting more things at the bottom 2

1 Introduction

This package allows the marking of (preliminary) versions of a document. This is done using the command \PrelimText, whose expansion is added below the footer of every page of a document (look at the bottom of this page for an example).

2 Usage

Simply using this package via \usepackage{prelim2e} produces a text in the form of “Preliminary version – March 6, 1996”.

\PrelimText The text is produced by the commands \PrelimText and \PrelimWords, which

\PrelimWords can be changed via \renewcommand* or by using options at the \usepackage

command (see section 3).

The footer for this documentation was produced by changing \PrelimWords to

\renewcommand*{\PrelimWords}{%
   \textsf{prelim2e} package --
   Version \fileversion{} --
   Documentation \LaTeX{}ed%
   }

1

∗The version number of this file is v1.10, last revised 1996/01/01.

prelim2e package – Version v1.10 – Documentation L^{A}TEXed – March 6, 1996 – 12:23

3 Options

The package has the following options: draft If this option is used a text appears below the normal pagebody. It is the

default. final If this option is used prelim2e produces no text. english This sets the text to “Preliminary version”. It is the default. german This sets the text to “Vorläufige Version”. It does not use the german or

babel package. time This adds the time at the beginning of the current L^{A}TEX to the text. The

time is produced using the scrtime package. All other options are passed to scrtime if the time option is selected.

4 Required packages

The package requires the following packages: everyshi It is used to implement the setting of the text below the normal pagebody. scrtime It is used to typeset the current time and is needed if the time option is

selected.

5 Putting more things at the bottom

Another package you might like to use with prelim2e is vrsion. This allows the definition and maintenance of a version number like 3.14 within L^{A}TEX, which can be put at the bottom of every page using prelime2e.

prelim2e package – Version v1.10 – Documentation L^{A}TEXed – March 6, 1996 – 12:23